GNU Info User’s Guide

For GNU Info version 2.11

Brian J. Fox (bfox@ai.mit.edu)

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993 Free Software Foundation

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the sections entitled “Copying” and “GNU General Public License” are included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

1 What is Info?

This file documents GNU Info, a program for viewing the on-line formatted versions of Texinfo files, version 2.11.

Info is a program which is used to view Info files on an ASCII terminal. Info files are the result of processing Texinfo files with the program makeinfo or with one of the Emacs commands, such as M-x texinfo-format-buffer. Texinfo itself is a documentation system that uses a single source file to produce both on-line information and printed output. You can typeset and print the files that you read in Info.

2 Command Line Options

GNU Info accepts several options to control the initial node being viewed, and to specify which directories to search for Info files. Here is a template showing an invocation of GNU Info from the shell:

info [--option-name option-value] menu-item…

The following option-names are available when invoking Info from the shell:

--directory directory-path -d directory-path

Add directory-path to the list of directory paths searched when Info needs to find a file. You may issue --directory multiple times; once for each directory which contains Info files. Alternatively, you may specify a value for the environment variable INFOPATH; if --directory is not given, the value of INFOPATH is used. The value of INFOPATH is a colon separated list of directory names. If you do not supply INFOPATH or -- directory-path, Info uses a default path.

--file filename -f filename

Specify a particular Info file to visit. By default, Info visits the file dir; if you use this option, Info will start with (filename)Top as the first file and node.

--node nodename -n nodename

Specify a particular node to visit in the initial file that Info loads. This is especially useful in conjunction with --file^{1}. You may specify --node multiple times; for an interactive Info, each nodename is visited in its own window, for a non-interactive Info (such as when --output is given) each nodename is processed sequentially.

--output filename

-o filename Specify filename as the name of a file to which to direct output. Each node that Info visits will be output to filename instead of interactively viewed. A value of - for filename specifies the standard output.

1 Of course, you can specify both the file and node in a --node command; but don’t forget to escape the open and close parentheses from the shell as in: info --node "(emacs)Buffers"

--subnodes

This option only has meaning when given in conjunction with --output. It means to recursively output the nodes appearing in the menus of each node being output. Menu items which resolve to external Info files are not output, and neither are menu items which are members of an index. Each node is only output once.

--help

-h Produces a relatively brief description of the available Info options.

--version

Prints the version information of Info and exits.

menu-item Info treats its remaining arguments as the names of menu items. The first argument is a menu item in the initial node visited, while the second argument is a menu item in the first argument’s node. You can easily move to the node of your choice by specifying the menu names which describe the path to that node. For example,

     info emacs buffers

first selects the menu item ‘Emacs’ in the node ‘(dir)Top’, and then selects the menu item ‘Buffers’ in the node ‘(emacs)Top’.

3 Moving the Cursor

Many people find that reading screens of text page by page is made easier when one is able to indicate particular pieces of text with some kind of pointing device. Since this is the case, GNU Info (both the Emacs and standalone versions) have several commands which allow you to move the cursor about the screen. The notation used in this manual to describe keystrokes is identical to the notation used within the Emacs manual, and the GNU Readline manual. See section “Character Conventions” in the GNU Emacs Manual, if you are unfamiliar with the notation.

The following table lists the basic cursor movement commands in Info. Each entry consists of the key sequence you should type to execute the cursor movement, the M-x^{1} command name (displayed in parentheses), and a short description of what the command does. All of the cursor motion commands can take an numeric argument (see Chapter 10 [Miscellaneous Commands], page 20), to find out how to supply them. With a numeric argument, the motion commands are simply executed that many times; for example, a numeric argument of 4 given to next-line causes the cursor to move down 4 lines. With a negative numeric argument, the motion is reversed; an argument of -4 given to the next-line command would cause the cursor to move up 4 lines.

C-n (next-line)

Move the cursor down to the next line.

C-p (prev-line)

Move the cursor up to the previous line.

C-a (beginning-of-line)

Move the cursor to the start of the current line.

C-e (end-of-line)

Move the cursor to the end of the current line.

C-f (forward-char)

Move the cursor forward a character.

C-b (backward-char)

Move the cursor backward a character.

M-f (forward-word)

Move the cursor forward a word.

M-b (backward-word)

Move the cursor backward a word.

1 M-x is also a command; it invokes execute-extended-command. See section “Executing an extended command” in the GNU Emacs Manual, for more detailed information.

M-< (beginning-of-node)

b Move the cursor to the start of the current node.

M-> (end-of-node)

Move the cursor to the end of the current node.

M-r (move-to-window-line)

Move the cursor to a specific line of the window. Without a numeric argument, M-r moves the cursor to the start of the line in the center of the window. With a numeric argument of n, M-r moves the cursor to the start of the nth line in the window.

4 Moving Text Within a Window

Sometimes you are looking at a screenful of text, and only part of the current paragraph you are reading is visible on the screen. The commands detailed in this section are used to shift which part of the current node is visible on the screen.

SPC (scroll-forward)

C-v Shift the text in this window up. That is, show more of the node which is currently below the bottom of the window. With a numeric argument, show that many more lines at the bottom of the window; a numeric argument of 4 would shift all of the text in the window up 4 lines (discarding the top 4 lines), and show you four new lines at the bottom of the window. Without a numeric argument, SPC takes the bottom two lines of the window and places them at the top of the window, redisplaying almost a completely new screenful of lines.

DEL (scroll-backward)

M-v Shift the text in this window down. The inverse of scroll-forward.

The scroll-forward and scroll-backward commands can also move forward and backward through the node structure of the file. If you press SPC while viewing the end of a node, or DEL while viewing the beginning of a node, what happens is controlled by the variable scroll-behavior. See Chapter 11 [Variables], page 22, for more information.

C-l (redraw-display)

Redraw the display from scratch, or shift the line containing the cursor to a specified location. With no numeric argument, ‘C-l’ clears the screen, and then redraws its entire contents. Given a numeric argument of n, the line containing the cursor is shifted so that it is on the nth line of the window.

C-x w (toggle-wrap)

Toggles the state of line wrapping in the current window. Normally, lines which are longer than the screen width wrap, i.e., they are continued on the next line. Lines which wrap have a ‘\’ appearing in the rightmost column of the screen. You can cause such lines to be terminated at the rightmost column by changing the state of line wrapping in the window with C-x w. When a line which needs more space than one screen width to display is displayed, a ‘$’ appears in the rightmost column of the screen, and the remainder of the line is invisible.

5 Selecting a New Node

This section details the numerous Info commands which select a new node to view in the current window.

The most basic node commands are ‘n’, ‘p’, ‘u’, and ‘l’.

When you are viewing a node, the top line of the node contains some Info pointers which describe where the next, previous, and up nodes are. Info uses this line to move about the node structure of the file when you use the following commands:

n (next-node)

Select the ‘Next’ node.

p (prev-node)

Select the ‘Prev’ node.

u (up-node)

Select the ‘Up’ node.

You can easily select a node that you have already viewed in this window by using the ‘l’ command – this name stands for "last", and actually moves through the list of already visited nodes for this window. ‘l’ with a negative numeric argument moves forward through the history of nodes for this window, so you can quickly step between two adjacent (in viewing history) nodes.

l (history-node)

Select the most recently selected node in this window.

Two additional commands make it easy to select the most commonly selected nodes; they are ‘t’ and ‘d’.

t (top-node)

Select the node ‘Top’ in the current Info file.

d (dir-node)

Select the directory node (i.e., the node ‘(dir)’).

Here are some other commands which immediately result in the selection of a different node in the current window:

< (first-node)

Selects the first node which appears in this file. This node is most often ‘Top’, but it does not have to be.

> (last-node)

Select the last node which appears in this file.

] (global-next-node)

Move forward or down through node structure. If the node that you are currently viewing has a ‘Next’ pointer, that node is selected. Otherwise, if this node has a menu, the first menu item is selected. If there is no ‘Next’ and no menu, the same process is tried with the ‘Up’ node of this node.

[ (global-prev-node)

Move backward or up through node structure. If the node that you are currently viewing has a ‘Prev’ pointer, that node is selected. Otherwise, if the node has an ‘Up’ pointer, that node is selected, and if it has a menu, the last item in the menu is selected.

You can get the same behavior as global-next-node and global-prev-node while simply scrolling through the file with SPC and DEL; See Chapter 11 [Variables], page 22, for more information.

g (goto-node)

Read the name of a node and select it. No completion is done while reading the node name, since the desired node may reside in a separate file. The node must be typed exactly as it appears in the Info file. A file name may be included as with any node specification, for example

     g(emacs)Buffers

finds the node ‘Buffers’ in the Info file ‘emacs’.

C-x k (kill-node)

Kill a node. The node name is prompted for in the echo area, with a default of the current node. Killing a node means that Info tries hard to forget about it, removing it from the list of history nodes kept for the window where that node is found. Another node is selected in the window which contained the killed node.

C-x C-f (view-file)

Read the name of a file and selects the entire file. The command

C-x C-f filename is equivalent to typing

g(filename)*

C-x C-b (list-visited-nodes)

Make a window containing a menu of all of the currently visited nodes. This window becomes the selected window, and you may use the standard Info commands within it.

C-x b (select-visited-node)

Select a node which has been previously visited in a visible window. This is similar to ‘C-x C-b’ followed by ‘m’, but no window is created.

6 Searching an Info File

GNU Info allows you to search for a sequence of characters throughout an entire Info file, search through the indices of an Info file, or find areas within an Info file which discuss a particular topic.

s (search)

Read a string in the echo area and search for it.

C-s (isearch-forward)

Interactively search forward through the Info file for a string as you type it.

C-r (isearch-backward)

Interactively search backward through the Info file for a string as you type it.

i (index-search)

Look up a string in the indices for this Info file, and select a node where the found index entry points to.

, (next-index-match)

Move to the node containing the next matching index item from the last ‘i’ command.

The most basic searching command is ‘s’ (search). The ‘s’ command prompts you for a string in the echo area, and then searches the remainder of the Info file for an occurrence of that string. If the string is found, the node containing it is selected, and the cursor is left positioned at the start of the found string. Subsequent ‘s’ commands show you the default search string within ‘[’ and ‘]’; pressing RET instead of typing a new string will use the default search string.

Incremental searching is similar to basic searching, but the string is looked up while you are typing it, instead of waiting until the entire search string has been specified.

7 Selecting Cross References

We have already discussed the ‘Next’, ‘Prev’, and ‘Up’ pointers which appear at the top of a node. In addition to these pointers, a node may contain other pointers which refer you to a different node, perhaps in another Info file. Such pointers are called cross references, or xrefs for short.

7.1 Parts of an Xref

Cross references have two major parts: the first part is called the label; it is the name that you can use to refer to the cross reference, and the second is the target; it is the full name of the node that the cross reference points to.

The target is separated from the label by a colon ‘:’; first the label appears, and then the target. For example, in the sample menu cross reference below, the single colon separates the label from the target.

     * Foo Label: Foo Target.        More information about Foo.

Note the ‘.’ which ends the name of the target. The ‘.’ is not part of the target; it serves only to let Info know where the target name ends.

A shorthand way of specifying references allows two adjacent colons to stand for a target name which is the same as the label name:

     * Foo Commands::                Commands pertaining to Foo.

In the above example, the name of the target is the same as the name of the label, in this case

Foo Commands.

You will normally see two types of cross reference while viewing nodes: menu references, and note references. Menu references appear within a node’s menu; they begin with a ‘*’ at the beginning of a line, and continue with a label, a target, and a comment which describes what the contents of the node pointed to contains.

Note references appear within the body of the node text; they begin with *Note, and continue with a label and a target.

Like ‘Next’, ‘Prev’, and ‘Up’ pointers, cross references can point to any valid node. They are used to refer you to a place where more detailed information can be found on a particular subject. Here is a cross reference which points to a node within the Texinfo documentation: See section “Writing an Xref” in the Texinfo Manual, for more information on creating your own texinfo cross references.

7.2 Selecting Xrefs

The following table lists the Info commands which operate on menu items.

1 (menu-digit)

2 … 9 Within an Info window, pressing a single digit, (such as ‘1’), selects that menu item, and places its node in the current window. For convenience, there is one exception; pressing ‘0’ selects the last item in the node’s menu.

0 (last-menu-item)

Select the last item in the current node’s menu.

m (menu-item)

Reads the name of a menu item in the echo area and selects its node. Completion is available while reading the menu label.

M-x find-menu

Move the cursor to the start of this node’s menu.

This table lists the Info commands which operate on note cross references.

f (xref-item)

r Reads the name of a note cross reference in the echo area and selects its node. Completion is available while reading the cross reference label.

Finally, the next few commands operate on menu or note references alike:

TAB (move-to-next-xref)

Move the cursor to the start of the next nearest menu item or note reference in this node. You can then use RET (select-reference-this-line) to select the menu or note reference.

M-TAB (move-to-prev-xref)

Move the cursor the start of the nearest previous menu item or note reference in this node.

RET (select-reference-this-line)

Select the menu item or note reference appearing on this line.

8 Manipulating Multiple Windows

A window is a place to show the text of a node. Windows have a view area where the text of the node is displayed, and an associated mode line, which briefly describes the node being viewed.

GNU Info supports multiple windows appearing in a single screen; each window is separated from the next by its modeline. At any time, there is only one active window, that is, the window in which the cursor appears. There are commands available for creating windows, changing the size of windows, selecting which window is active, and for deleting windows.

8.1 The Mode Line

A mode line is a line of inverse video which appears at the bottom of an Info window. It describes the contents of the window just above it; this information includes the name of the file and node appearing in that window, the number of screen lines it takes to display the node, and the percentage of text that is above the top of the window. It can also tell you if the indirect tags table for this Info file needs to be updated, and whether or not the Info file was compressed when stored on disk.

Here is a sample mode line for a window containing an uncompressed file named ‘dir’, showing the node ‘Top’.

     -----Info: (dir)Top, 40 lines --Top---------------------------------------
                 ^^^^^^^^
               (file)Node #lines    where

When a node comes from a file which is compressed on disk, this is indicated in the mode line with two small ‘z’’s. In addition, if the Info file containing the node has been split into subfiles, the name of the subfile containing the node appears in the modeline as well:

     --zz-Info: (emacs)Top, 291 lines --Top-- Subfile: emacs-1.Z---------------

When Info makes a node internally, such that there is no corresponding info file on disk, the name of the node is surrounded by asterisks (‘*’). The name itself tells you what the contents of the window are; the sample mode line below shows an internally constructed node showing possible completions:

     -----Info: *Completions*, 7 lines --All-----------------------------------

8.2 Window Commands

It can be convenient to view more than one node at a time. To allow this, Info can display more than one window. Each window has its own mode line (see Section 8.1 [The Mode Line], page 14) and history of nodes viewed in that window (see Chapter 5 [history-node], page 7).

C-x o (next-window)

Select the next window on the screen. Note that the echo area can only be selected if it is already in use, and you have left it temporarily. Normally, ‘C-x o’ simply moves the cursor into the next window on the screen, or if you are already within the last window, into the first window on the screen. Given a numeric argument, ‘C-x o’ moves over that many windows. A negative argument causes ‘C-x o’ to select the previous window on the screen.

M-x prev-window

Select the previous window on the screen. This is identical to ‘C-x o’ with a negative argument.

C-x 2 (split-window)

Split the current window into two windows, both showing the same node. Each window is one half the size of the original window, and the cursor remains in the original window. The variable automatic-tiling can cause all of the windows on the screen to be resized for you automatically, please see Chapter 11 [automatic-tiling], page 22 for more information.

C-x 0 (delete-window)

Delete the current window from the screen. If you have made too many windows and your screen appears cluttered, this is the way to get rid of some of them.

C-x 1 (keep-one-window)

Delete all of the windows excepting the current one.

ESC C-v (scroll-other-window)

Scroll the other window, in the same fashion that ‘C-v’ might scroll the current window. Given a negative argument, scroll the "other" window backward.

C-x ^(grow-window)

Grow (or shrink) the current window. Given a numeric argument, grow the current window that many lines; with a negative numeric argument, shrink the window instead.

C-x t (tile-windows)

Divide the available screen space among all of the visible windows. Each window is given an equal portion of the screen in which to display its contents. The variable automatic-tiling can cause tile-windows to be called when a window is created or deleted. See Chapter 11 [automatic-tiling], page 22.

8.3 The Echo Area

The echo area is a one line window which appears at the bottom of the screen. It is used to display informative or error messages, and to read lines of input from you when that is necessary. Almost all of the commands available in the echo area are identical to their Emacs counterparts, so please refer to that documentation for greater depth of discussion on the concepts of editing a line of text. The following table briefly lists the commands that are available while input is being read in the echo area:

C-f (echo-area-forward)

Move forward a character.

C-b (echo-area-backward)

Move backward a character.

C-a (echo-area-beg-of-line)

Move to the start of the input line.

C-e (echo-area-end-of-line)

Move to the end of the input line.

M-f (echo-area-forward-word)

Move forward a word.

M-b (echo-area-backward-word)

Move backward a word.

C-d (echo-area-delete)

Delete the character under the cursor.

DEL (echo-area-rubout)

Delete the character behind the cursor.

C-g (echo-area-abort)

Cancel or quit the current operation. If completion is being read, ‘C-g’ discards the text of the input line which does not match any completion. If the input line is empty, ‘C-g’ aborts the calling function.

RET (echo-area-newline)

Accept (or forces completion of) the current input line.

C-q (echo-area-quoted-insert)

Insert the next character verbatim. This is how you can insert control characters into a search string, for example.

printing character (echo-area-insert)

Insert the character.

M-TAB (echo-area-tab-insert)

Insert a TAB character.

C-t (echo-area-transpose-chars)

Transpose the characters at the cursor.

The next group of commands deal with killing, and yanking text. For an in depth discussion of killing and yanking, see section “Killing and Deleting” in the GNU Emacs Manual

M-d (echo-area-kill-word)

Kill the word following the cursor.

M-DEL (echo-area-backward-kill-word)

Kill the word preceding the cursor.

C-k (echo-area-kill-line)

Kill the text from the cursor to the end of the line.

C-x DEL (echo-area-backward-kill-line)

Kill the text from the cursor to the beginning of the line.

C-y (echo-area-yank)

Yank back the contents of the last kill.

M-y (echo-area-yank-pop)

Yank back a previous kill, removing the last yanked text first.

Sometimes when reading input in the echo area, the command that needed input will only accept one of a list of several choices. The choices represent the possible completions, and you must respond with one of them. Since there are a limited number of responses you can make, Info allows you to abbreviate what you type, only typing as much of the response as is necessary to uniquely identify it. In addition, you can request Info to fill in as much of the response as is possible; this is called completion.

The following commands are available when completing in the echo area:

TAB (echo-area-complete)

SPC Insert as much of a completion as is possible.

? (echo-area-possible-completions)

Display a window containing a list of the possible completions of what you have typed so far. For example, if the available choices are:

     bar foliate food forget

and you have typed an ‘f’, followed by ‘?’, the possible completions would contain:

     foliate food forget

i.e., all of the choices which begin with ‘f’. Pressing SPC or TAB would result in ‘fo’ appearing in the echo area, since all of the choices which begin with ‘f’ continue with ‘o’. Now, typing ‘l’ followed by ‘TAB’ results in ‘foliate’ appearing in the echo area, since that is the only choice which begins with ‘fol’.

ESC C-v (echo-area-scroll-completions-window)

Scroll the completions window, if that is visible, or the "other" window if not.

9 Printing Out Nodes

You may wish to print out the contents of a node as a quick reference document for later use. Info provides you with a command for doing this. In general, we recommend that you use TEX to format the document and print sections of it, by running tex on the Texinfo source file.

M-x print-node

Pipe the contents of the current node through the command in the environment variable INFO_PRINT_COMMAND. If the variable does not exist, the node is simply piped to lpr.

10 Miscellaneous Commands

GNU Info contains several commands which self-document GNU Info:

M-x describe-command

Read the name of an Info command in the echo area and then display a brief description of what that command does.

M-x describe-key

Read a key sequence in the echo area, and then display the name and documentation of the Info command that the key sequence invokes.

M-x describe-variable

Read the name of a variable in the echo area and then display a brief description of what the variable affects.

M-x where-is

Read the name of an Info command in the echo area, and then display a key sequence which can be typed in order to invoke that command.

C-h (get-help-window)

? Create (or Move into) the window displaying *Help*, and place a node containing a quick reference card into it. This window displays the most concise information about GNU Info available.

h (get-info-help-node)

Try hard to visit the node (info)Help. The Info file ‘info.texi’ distributed with GNU Info contains this node. Of course, the file must first be processed with makeinfo, and then placed into the location of your Info directory.

Here are the commands for creating a numeric argument:

C-u (universal-argument)

Start (or multiply by 4) the current numeric argument. ‘C-u’ is a good way to give a small numeric argument to cursor movement or scrolling commands; ‘C-u C-v’ scrolls the screen 4 lines, while ‘C-u C-u C-n’ moves the cursor down 16 lines.

M-1 (add-digit-to-numeric-arg)
M-2 . . . M-9

Add the digit value of the invoking key to the current numeric argument. Once Info is reading a numeric argument, you may just type the digits of the argument, without the Meta prefix. For example, you might give ‘C-l’ a numeric argument of 32 by typing:

               C-u 3 2 C-l

or

               M-3 2 C-l

‘C-g’ is used to abort the reading of a multi-character key sequence, to cancel lengthy operations (such as multi-file searches) and to cancel reading input in the echo area.

C-g (abort-key)

Cancel current operation.

The ‘q’ command of Info simply quits running Info.

q (quit) Exit GNU Info.

If the operating system tells GNU Info that the screen is 60 lines tall, and it is actually only 40 lines tall, here is a way to tell Info that the operating system is correct.

M-x set-screen-height

Read a height value in the echo area and set the height of the displayed screen to that value.

Finally, Info provides a convenient way to display footnotes which might be associated with the current node that you are viewing:

ESC C-f (show-footnotes)

Show the footnotes (if any) associated with the current node in another window. You can have Info automatically display the footnotes associated with a node when the node is selected by setting the variable automatic-footnotes. See Chapter 11

          [automatic-footnotes], page 22.

11 Manipulating Variables

GNU Info contains several variables whose values are looked at by various Info commands. You can change the values of these variables, and thus change the behavior of Info to more closely match your environment and Info file reading manner.

M-x set-variable

Read the name of a variable, and the value for it, in the echo area and then set the variable to that value. Completion is available when reading the variable name; often, completion is available when reading the value to give to the variable, but that depends on the variable itself. If a variable does not supply multiple choices to complete over, it expects a numeric value.

M-x describe-variable

Read the name of a variable in the echo area and then display a brief description of what the variable affects.

Here is a list of the variables that you can set in Info.

automatic-footnotes

When set to On, footnotes appear and disappear automatically. This variable is On by default. When a node is selected, a window containing the footnotes which appear in that node is created, and the footnotes are displayed within the new window. The window that Info creates to contain the footnotes is called ‘*Footnotes*’. If a node is selected which contains no footnotes, and a ‘*Footnotes*’ window is on the screen, the ‘*Footnotes*’ window is deleted. Footnote windows created in this fashion are not automatically tiled so that they can use as little of the display as is possible.

automatic-tiling

When set to On, creating or deleting a window resizes other windows. This variable is Off by default. Normally, typing ‘C-x 2’ divides the current window into two equal parts. When automatic-tiling is set to On, all of the windows are resized automatically, keeping an equal number of lines visible in each window. There are exceptions to the automatic tiling; specifically, the windows ‘*Completions*’ and ‘*Footnotes*’ are not resized through automatic tiling; they remain their original size.

visible-bell

When set to On, GNU Info attempts to flash the screen instead of ringing the bell. This variable is Off by default. Of course, Info can only flash the screen if the terminal allows it; in the case that the terminal does not allow it, the setting of this variable has no effect. However, you can make Info perform quietly by setting the errors-ring-bell variable to Off.

errors-ring-bell

When set to On, errors cause the bell to ring. The default setting of this variable is On.

gc-compressed-files

When set to On, Info garbage collects files which had to be uncompressed. The default value of this variable is Off. Whenever a node is visited in Info, the Info file containing that node is read into core, and Info reads information about the tags and nodes contained in that file. Once the tags information is read by Info, it is never forgotten. However, the actual text of the nodes does not need to remain in core unless a particular Info window needs it. For non-compressed files, the text of the nodes does not remain in core when it is no longer in use. But de-compressing a file can be a time consuming operation, and so Info tries hard not to do it twice. gc-compressed-files tells Info it is okay to garbage collect the text of the nodes of a file which was compressed on disk.

show-index-match

When set to On, the portion of the matched search string is highlighted in the message which explains where the matched search string was found. The default value of this variable is On. When Info displays the location where an index match was found, (see Chapter 6 [next-index-match], page 10), the portion of the string that you had typed is highlighted by displaying it in the inverse case from its surrounding characters.

scroll-behavior

Control what happens when forward scrolling is requested at the end of a node, or when backward scrolling is requested at the beginning of a node. The default value for this variable is Continuous. There are three possible values for this variable:

Continuous

Try to get the first item in this node’s menu, or failing that, the ‘Next’ node, or failing that, the ‘Next’ of the ‘Up’. This behavior is identical to using

          the ‘]’ (global-next-node) and ‘[’ (global-prev-node) commands.

Next Only Only try to get the ‘Next’ node.

Page Only Simply give up, changing nothing. If scroll-behavior is Page Only, no scrolling command can change the node that is being viewed.

scroll-step

The number of lines to scroll when the cursor moves out of the window. Scrolling happens automatically if the cursor has moved out of the visible portion of the node text when it is time to display. Usually the scrolling is done so as to put the cursor on the center line of the current window. However, if the variable scroll-step has a nonzero value, Info attempts to scroll the node text by that many lines; if that is enough to bring the cursor back into the window, that is what is done. The default value of this variable is 0, thus placing the cursor (and the text it is attached to) in the center of the window. Setting this variable to 1 causes a kind of "smooth scrolling" which some people prefer.

ISO-Latin

When set to On, Info accepts and displays ISO Latin characters. By default, Info assumes an ASCII character set. ISO-Latin tells Info that it is running in an environment where the European standard character set is in use, and allows you to input such characters to Info, as well as display them.

Appendix A: Global Index 25

Appendix A Global Index

,

,. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

-

--subnodes, command line option . . . . . . . . . . . . . . . . . . . 3

?

?, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ?, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

[

[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

]

] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

0

0, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

1

1 … 9, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

A

abort-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 add-digit-to-numeric-arg . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 arguments, command line . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 automatic-footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 automatic-tiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

B

b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 backward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 backward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 beginning-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 beginning-of-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

C

C-a, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-a, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C-b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-b, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C-d, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C-e, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-e, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C-f, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-f, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C-g, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 C-g, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C-h. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 C-k, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 C-l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-q, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 C-t, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 C-u. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-x ^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x 0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-x C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-x C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-x DEL, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 17 C-x k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-x o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-x t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-y, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 cancelling the current operation . . . . . . . . . . . . . . . . . . . . 21 cancelling typeahead . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 command line options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 commands, describing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 cursor, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 DEL, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 DEL, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 delete-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 describe-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 describe-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 describe-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 dir-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 directory path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

E

echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-backward-kill-line . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-backward-kill-word. . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-backward-word . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-beg-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-complete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-delete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-end-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-insert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-kill-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-kill-word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-possible-completions . . . . . . . . . . . . . . . . . . . . . 18 echo-area-quoted-insert . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-rubout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-scroll-completions-window. . . . . . . . . . . . . . . . 18 echo-area-tab-insert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-transpose-chars . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-yank. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 echo-area-yank-pop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 end-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 end-of-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 errors-ring-bell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ESC C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ESC C-v, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . 15 ESC C-v, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 18

F

f. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 file, outputting to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 find-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 first-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 footnotes, displaying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 forward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 functions, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 gc-compressed-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 get-help-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 get-info-help-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 global-next-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 global-prev-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 goto-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 grow-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 history-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

I

i. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 index-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Info file, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 INFO PRINT COMMAND, environment variable . . 19 isearch-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 isearch-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ISO Latin characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ISO-Latin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

K

keep-one-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 keys, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 kill-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

L

l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 last-menu-item. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 last-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 list-visited-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-1 … M-9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 M-b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 M-b, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 M-d, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-DEL, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-f, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 M-f, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 M-r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-TAB, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 M-TAB, in the echo area. . . . . . . . . . . . . . . . . . . . . . . . . . . 17 M-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 M-y, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 menu, following . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 menu-digit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 move-to-next-xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 move-to-prev-xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 move-to-window-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 next-index-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 next-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 next-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 next-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 node, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 nodes, selection of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 numeric arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

O

outputting to a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 prev-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 prev-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 prev-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 print-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 printing characters, in the echo area. . . . . . . . . . . . . . . . 17

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 quitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 redraw-display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 RET, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 RET, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 screen, changing the height of . . . . . . . . . . . . . . . . . . . . . . 21 scroll-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 scroll-behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 scroll-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 scroll-other-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 scroll-step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 scrolling through node structure. . . . . . . . . . . . . . . . . . . . . 6 search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 select-reference-this-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 select-visited-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 set-screen-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 set-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 show-footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 show-index-match. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 SPC, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 SPC, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 split-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

T

t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 TAB, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 TAB, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 tile-windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 tiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 toggle-wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 top-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

U

u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 universal-argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 up-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

V

variables, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 variables, setting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 version information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 view-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 visible-bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

W

where-is . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 windows, creating. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 windows, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 windows, manipulating. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 windows, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

X

xref-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Table of Contents

1 What is Info?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Command Line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Moving the Cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4 Moving Text Within a Window . . . . . . . . . . . . . . . . . . . . 6 5 Selecting a New Node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6 Searching an Info File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 7 Selecting Cross References. . . . . . . . . . . . . . . . . . . . . . . . 11

7.1 Parts of an Xref. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 7.2 Selecting Xrefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

8 Manipulating Multiple Windows . . . . . . . . . . . . . . . . . 14

8.1 The Mode Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.2 Window Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 8.3 The Echo Area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

9 Printing Out Nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 10 Miscellaneous Commands. . . . . . . . . . . . . . . . . . . . . . . . 20 11 Manipulating Variables . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Appendix A Global Index . . . . . . . . . . . . . . . . . . . . . . . . . . 25